OpenStudy (anonymous):

Factor: x2 - 13x + 36

OpenStudy (anonymous):

the way to do this is to ask : what two factors of 36 add up to -13

OpenStudy (anonymous):

(x + b)(x + d) = 0 is what you want to set up then as @thomasj said you want to figure out what two factors of 36 add up to -13. meaning what b*d = 36 and also b+d = -13

OpenStudy (precal):

 continue writing the factors, and you will see what you need